# R u s s i a n   A c a d e m y   o f   S c i e n c e s  S i b e r i a n    B r a n c h

### v37

@@ I. Kapovich
Yes, it is (E. Khukhro, Preprint, 2025, http://arxiv.org/abs/2505.20999).
@@ 19.53. Let $G$ be a group generated by elements $x,y,z$ such that
19.60. Let $R$ be a principal ideal domain, let $d$ be a positive integer, and let $X_{1},\,\ldots,\,X_{m}$ be members of $\mathrm{SL}(d,R)$ (or of $\mathrm{GL}(d,R)$). Is it decidable whether or not these $m$ matrices freely generate a free group? If it is, design an effective algorithm for computing the answer. Peter M. Neumann
@@
19.62. Let $\mathfrak{A}=\{\mathfrak{A}_{r}\ |\ r\in\Phi\}$ be an elementary carpet of type $\Phi$ of rank $l\geqslant 2$ (see 7.28). For $p\in\Phi$, define a set of additive subgroups $\mathfrak{B}_{p}=\sum C_{ij,rs}\mathfrak{A}_{r}^{i}\mathfrak{A}_{s}^{j}$, where the sum is taken over all natural numbers $i$, $j$ and roots $r,s\in\Phi$ such that $ir+js=p$. It is known that the set $\mathfrak{B}=\{\mathfrak{B}_{p}\ |\ p\in\Phi\}$ is a carpet called the derived carpet of $\mathfrak{A}$. It is also known that for $\Phi=A_{l}$ the set $\mathfrak{B}$ is a closed (admissible) carpet, which means that its carpet subgroup does not contain new root elements. Is every derived carpet of type $\Phi$ over a commutative ring closed (admissible)? Ya. N. Nuzhin
@@ I. Ponomarenko
Yes, it does (S. Corson, Preprint, 2025, https://arxiv.org/abs/2505.07330).
@@ V. Shpilrain
Yes, it is true (A.-M. Liu, W. Guo, I. N. Safonova, A. N. Skiba, to appear in J. Algebra, 2021); another solution using CFSG is in (S. F. Kamornikov, V. N. Tyutyanov, Trudy Inst. Mat. Mekh. UrO RAN, 27, no. 1 (2021), 98–102 (Russian);  X. Yi, S. F. Kamornikov, V. N. Tyutyanov, Probl. Physics, Math. Techn., 46, no. 1 (2021), 50–53).
@@
Yes, it is true (A.-M. Liu, W. Guo, I. N. Safonova, A. N. Skiba, to appear in J. Algebra, 2021); another solution using CFSG is in (S. F. Kamornikov, V. N. Tyutyanov, Trudy Inst. Mat. Mekh. UrO RAN, 27, no. 1 (2021), 98–102 (Russian);  X. Yi, S. F. Kamornikov, V. N. Tyutyanov, Probl. Physics, Math. Techn., 46, no. 1 (2021), 50–53).
@@
Yes, it is (L. S. Kazarin, V. N. Tyutyanov, Preprint, 2023).
@@
19.105. Is it true that fixed subgroups of endomorphisms of $F$ are inert?
It is easy to see that retracts of free groups are compressed, and that they are inert if and only if fixed subgroups of endomorphisms of free groups are inert. Therefore the question is equivalent to the following: are retracts of free groups inert? E. Ventura
@@ 20.2. Are there any nonabelian simple groups of Lie type which are totally $3$-closed?
20.3. For each $k$, each group of order $k$ is totally $k$-closed, while the alternating group $A_{k}$ is not totally k2k-2-closed. Is there a fixed integer $k$ such that all finite simple groups which are not alternating groups are totally $k$-closed?
@@
a) Is the same true for a definably simple locally finite group? A group is said to be definably simple if it does not contain proper definable normal subgroups.
b) The same question for a definably simple locally finite group with bounded derived lengths of its solvable subgroups. A. V. Borovik
@@
(ao) What is the exact value of  $\liminf_{n\ \mathrm{odd}}\gamma(S_{n})/n$? Comment of 2025: It belongs to [1/12, 1/6][1/12,\,1/6] for odd (S. Eberhard, C. Mellon, Preprint, 2024, https://arxiv.org/abs/2410.06999).
@@
(d) What are the exact values of  $\limsup_{n\ \mathrm{odd}}\gamma(A_{n})/n$   and   $\limsup_{n\ \mathrm{even}}\gamma(A_{n})/n$? D. Bubboloni, C. E. Praeger, P. Spiga
@@
20.20. Let $E(l,m,n)$ be a group on generators $a,b$ with the presentation $\langle a,b\mid a^{l}=\nobreak 1,\;(ab)^{m}=b^{n}\rangle$, where $l,m,n$ satisfy $1/l+1/m+1/n<1$. Are there any values for $l,m,n$ such that $E(l,m,n)$ is shortlex automatic?
@@ 20.24. Does the group $SL_{2}(\mathbb{Z}[\sqrt{2}])$ admit a faithful transitive amenable action?
M. H. Hooshmand
@@
The answer is known to be affirmative if $\lambda=\kappa$ (B. Kaya, M. Kuzucuoğlu, P. Longobardi, M. Maj, Preprint, 2024, https://arxiv.org/pdf/2409.00545.
B. Kaya, M. Kuzucuoğlu
@@
20.42. (Y. Cornullier, A. Mann, A. Thom). Does there exist a finitely generated residually finite group that is not (elementary) amenable and satisfies a nontrivial group law? S. Kionke
@@ E. Plotkin
The conjecture is proved when $m$ is a prime power (G. Qian, Arch. Math., 97 (2011), 99–103); when $m$ is square-free (I. M. Isaacs, Arch. Math., 97 (2011), 499–501); or when $G$ is solvable (G. Qian, Bull. London Math. Soc., 53 (2021) 820–824); or when $G$ is a symmetric or alternating group (E. Giannelli (J. Algebra Appl., 23, no. 9 (2024), article ID 2450144). G. Qian
20.79. Conjecture: Suppose that $G$ is a non-abelian simple group and $H$ is a finite group such that ${\rm Cod}(G)={\rm Cod}(H)$. Then $G\cong H$. G. Qian
@@ 20.81. Does there exist a class $\mathfrak{X}$ of finite groups satisfying the following conditions:
One can prove that there are no classes $\mathfrak{X}$ satisfying (1)–(3) that are closed with respect to extensions. D. O. Revin
@@
The structure of profinite CN-groups is described in (Israel J. Math., 235, no. 1 (2020), 325–347). P. Shumyatsky
@@ 20.97. Is a 2-group locally finite if the centralizer of every involution is locally finite?
It is known that in any group a 3-Engel element of odd order belongs to the locally nilpotent radical (E. Jabara, G. Traustason, Proc. Amer. Math. Soc., 147, no. 5 (2019), 1921–1927). G. Traustason
@@
If it can, then the answer to 20.117 is negative, but in the abstract case it is known that it can not. P. Zalesski
@@
Yes, there exist (V. S. Guba, Math. USSR–Izv., 29 (1987), 233–277).
@@
Yes, there exist (S. D. Brodskiĭ, Dep. no. 2214-80, VINITI, Moscow, 1980 (Russian)).
@@
Yes, there exists (L. A. Bokut’, Soviet Math. Dokl., 8 (1967), 901–904;  A. Bowtell, J. Algebra, 7 (1967), 126–139;  A. Klein, J. Algebra, 7 (1967), 100–125).
@@
Yes, there exists (G. Higman, Finitely presented infinite simple groups ((Notes on Pure Mathematics, no. $8)$ Dept. of Math., Inst. Adv. Studies Austral. National Univ., Canberra, 1974), where there is also a reference to R. Thompson.
@@
a) Yes, there exists (G. S. Makanin, Math. USSR–Izv., 21 (1982), 546–582; 25 (1985), 75–88).
@@ 1.45. Is the largest locally nilpotent normal subgroup relatively convex?
Yes, there exist (V. V. Bludov, Algebra and Logic, 13 (1974), 343–360).
@@ 1.61. Can any orderable group be embedded into an orderable group
a) Yes, there exist; b) Not necessarily (S. I. Adian, Math. USSR–Izv., 5 (1971), 475–484).
@@
Yes, there exist (E. M. Levich, A. I. Tokarenko, Siberian Math. J., 11 (1970), 1033–1034;  A. I. Tokarenko, Trudy Rizhsk. Algebr. Sem., Riga Univ., Riga, 1969, 280–281 (Russian)).
@@ 1.76. Does there exist a simple, locally nilpotent, locally compact, topological group?
Yes, there exists (V. S. Atabekyan, Dep. no. 5381-V86, VINITI, Moscow, 1986 (Russian)).
@@ 1.85. Is it true that the identity relations of a metabelian group have a finite basis?
Yes, they exist (S. V. Ivanov, Trans. Moscow Math. Soc., 1993, 217–249).

### v38

@@ I. Kapovich
Yes, it is (E. Khukhro, to appear in Bull. London Math. Soc., 2025, http://arxiv.org/abs/2505.20999).
@@ 19.53. Let $G$ be a group generated by elements $x,y,z$ such that
19.60. Let $R$ be a principal ideal domain, let $d$ be a positive integer, and let $X_{1},\,\ldots,\,X_{m}$ be members of $\mathrm{SL}(d,R)$ (or of $\mathrm{GL}(d,R)$). Is it decidable whether or not these $m$ matrices freely generate a free group? If it is, design an effective algorithm for computing the answer. P. M. Neumann
@@
19.62. Let $\mathfrak{A}=\{\mathfrak{A}_{r}\ |\ r\in\Phi\}$ be an elementary carpet of type $\Phi$ of rank $l\geqslant 2$ (see 7.28). For $p\in\Phi$, define a set of additive subgroups $\mathfrak{B}_{p}=\sum C_{ij,rs}\mathfrak{A}_{r}^{i}\mathfrak{A}_{s}^{j}$, where the sum is taken over all natural numbers $i$, $j$ and roots $r,s\in\Phi$ such that $ir+js=p$. It is known that the set $\mathfrak{B}=\{\mathfrak{B}_{p}\ |\ p\in\Phi\}$ is a carpet called the derived carpet of $\mathfrak{A}$. It is also known that for $\Phi=A_{l}$ the set $\mathfrak{B}$ is a closed (admissible) carpet, which means that its carpet subgroup does not contain new root elements. Is every derived carpet of type $\Phi$ over a commutative ring closed (admissible)?
Comment of 2025: An affirmative answer was obtained for $\Phi$ of type $B_{l}$, $C_{l}$, or $F_{4}$ when $\mbox{GCD}(p,2)=1$, and for $\Phi$ of type $G_{2}$ when $\mbox{GCD}(p,6)=1$ (Ya. N. Nuzhin, J. Siberian Fed. Univ. Ser. Math. Phys., 16, no. 6 (2023), 732–737). Ya. N. Nuzhin
@@
Yes, they are sufficient (Ya. N. Nuzhin, J. Siberian Fed. Univ. Ser. Math. Phys., 16, no. 6 (2023), 732–737; Ya. N. Nuzhin, to appear in Siberian Math. J.).
@@ I. Ponomarenko
Yes, it does (S. Corson, Monatsh. Math. (2025) https://doi.org/10.1007/s00605-025-02116-8.
@@ V. Shpilrain
Yes, it is true (A.-M. Liu, W. Guo, I. N. Safonova, A. N. Skiba, J. Algebra, 585 (2021), 280–293); another solution using CFSG is in (S. F. Kamornikov, V. N. Tyutyanov, Trudy Inst. Mat. Mekh. UrO RAN, 27, no. 1 (2021), 98–102 (Russian); X. Yi, S. F. Kamornikov, V. N. Tyutyanov, Probl. Physics, Math. Techn., 46, no. 1 (2021), 50–53).
@@
Yes, it is true (A.-M. Liu, W. Guo, I. N. Safonova, A. N. Skiba, J. Algebra, 585 (2021), 280–293); another solution using CFSG is in (S. F. Kamornikov, V. N. Tyutyanov, Trudy Inst. Mat. Mekh. UrO RAN, 27, no. 1 (2021), 98–102 (Russian); X. Yi, S. F. Kamornikov, V. N. Tyutyanov, Probl. Physics, Math. Techn., 46, no. 1 (2021), 50–53).
@@
It is determined: $Aut(A)\cong G_{2}\times SO(3)$, which follows from the result about derivations in (H. Petyt, Commun. Algebra, 47, no. 10 (2019), 4216–4223).
@@
Yes, it is (L. S. Kazarin, V. N. Tyutyanov, Preprint, 2023 (Russian),
https://kourovkanotebookorg.wordpress.com/wp-content/uploads/2025/09/d09ad0bed183d180d0bed0b2d0bad0b0-19.100-5.pdf).
@@
19.105. Is it true that the fixed subgroups of endomorphisms of $F$ are inert?
E. Ventura
@@ 20.2. Are there any nonabelian simple groups of Lie type which are totally $3$-closed?
20.3. For each $k$, each group of order $k$ is totally $k$-closed, while the alternating group $A_{k}$ is not totally (k2)(k-2)-closed. Is there a fixed integer $k$ such that all finite simple groups which are not alternating groups are totally $k$-closed?
@@
a) Is the same true for an infinite definably simple locally finite group? A group is said to be definably simple if it does not contain proper definable normal subgroups.
b) The same question for an infinite definably simple locally finite group with bounded derived lengths of its solvable subgroups. A. V. Borovik
@@
(ao) What is the exact value of $\liminf_{n\ \mathrm{odd}}\gamma(S_{n})/n$?
Comment of 2025: It belongs to [1/12, 1/6][1/12,\,1/6] for odd (S. Eberhard, C. Mellon, Preprint, 2024, https://arxiv.org/abs/2410.06999).
@@
(d) What are the exact values of $\limsup_{n\ \mathrm{odd}}\gamma(A_{n})/n$ and $\limsup_{n\ \mathrm{even}}\gamma(A_{n})/n$?
D. Bubboloni, C. E. Praeger, P. Spiga
@@
[…]
@@ 20.24. Does the group $SL_{2}(\mathbb{Z}[\sqrt{2}])$ admit a faithful transitive amenable action?
Comment of 2025: any minimal counterexample $G$ must be a simple group without prime-index subgroups (R. R. Bildanov, V. A. Goryachenko, A. V. Vasil’ev, Siberian Electron. Math. Rep., 17 (2020), 683–689; M. H. Hooshmand, Commun. Algebra, 49, no. 7 (2021), 2927–2933). M. H. Hooshmand
@@
The answer is known to be affirmative if $\lambda=\kappa$ (B. Kaya, M. Kuzucuoğlu, P. Longobardi, M. Maj, J. Algebra, 666 (2025), 840–849). B. Kaya, M. Kuzucuoğlu
@@
20.42. (Y. Cornulier, A. Mann, A. Thom). Does there exist a finitely generated residually finite group that is not (elementary) amenable and satisfies a nontrivial group law? S. Kionke
@@ E. Plotkin
The conjecture is proved when $m$ is a prime power (G. Qian, Arch. Math., 97 (2011), 99–103); when $m$ is square-free (I. M. Isaacs, Arch. Math., 97 (2011), 499–501); or when $G$ is solvable (G. Qian, Bull. London Math. Soc., 53 (2021) 820–824); or when $G$ is a symmetric or alternating group (E. Giannelli (J. Algebra Appl., 23, no. 9 (2024), article ID 2450144); or when $G$ is an almost simple group (S. Y. Madanha, Commun. Algebra, 51, no. 7 (2023), 3143–3151); or when $\mathbf{F}(G)=1$ (Z. Akhlaghi, E. Pacifici, L. Sanus, J. Algebra, 644 (2024), 428–441). G. Qian
20.79. Conjecture: Suppose that $G$ is a non-abelian simple group and $H$ is a finite group such that ${\rm Cod}(G)={\rm Cod}(H)$. Then $G\cong H$.
Comment of 2025: The conjecture is proved for $G$ of type $\mathrm{PSL}(2,q)$ (A. Bahri, Z. Akhlaghi, B. Khosravi, Bull. Austral. Math. Soc., 104, no. 2 (2021), 278–286); $\mathrm{PSL}(3,q)$ or $\mathrm{PSU}(3,q)$ (Y. Liu, Y. Yang, Results Math., 78, no. 1 (2023), article No. 7); a sporadic simple group (M. Dolorfino, L. Martin, Z. Slonim, Y. Sun, Y. Yang, Bull. Austral. Math. Soc., 109, no. 1 (2024), 57–66); an alternating group (M. Dolorfino, L. Martin, Z. Slonim, Y. Sun, Y. Yang, Bull. Austral. Math. Soc., 110, no. 1 (2024), 115–120); $^{2}F_{4}(q^{2})$ (Y. Yang, J. Group Theory, 27, no. 1 (2024), 141–155); $\mathrm{PSL}(n,q)$ or a simple exceptional group of Lie type (H. P. Tong Viet, Math. Nachr., 298, no. 4 (2025), 1356–1369). G. Qian
@@ 20.81. Does there exist a class $\mathfrak{X}$ of finite groups satisfying the following conditions:
It is known that there are no classes $\mathfrak{X}$ satisfying (1)–(3) that are closed with respect to extensions (D. O. Revin, Algebra i Analiz, 37, no. 1 (2025), 141–176 (Russian)). D. O. Revin
@@
Yes, it is true (L. Koch-Hyde, S. O’Connor, E. Olive, V. Shpilrain, Preprint, 2025, https://arxiv.org/abs/2505.00477).
@@
The structure of profinite CN-groups is described in (P. Shumyatsky, Israel J. Math., 235, no. 1 (2020), 325–347). P. Shumyatsky
@@
Comment of 2025: If $2\equiv\xi^{p^{n-1}}$ when $\xi$ is a primitive root modulo $p$, and if $A$ is strongly nilpotent of nilpotency index less than $p$, then the result follows by (A. Smoktunowicz, Adv. Math., 409, part B (2022), 108683). It is not known if the result follows when $A$ is not strongly nilpotent.
@@ 20.97. Is a 2-group locally finite if the centralizer of every involution is locally finite?
No, not necessarily (P. Neumann, Illinois J. Math., 30 (1986), 301–316). Moreover, there are finitely generated branch groups all of whose maximal subgroups are of finite index but not all of them are normal (M. Garciarena, M. Petschick, Preprint, 2024, https://arxiv.org/abs/2410.06783).
@@
It is known that in any group a 3-Engel element of odd order belongs to the locally nilpotent radical (E. Jabara, G. Traustason, Proc. Amer. Math. Soc., 147, no. 5 (2019), 1921–1927).
Comment of 2025: An element $a\in G$ is called a strong left $3$-Engel element if $\langle a,a^{g}\rangle$ is nilpotent of class at most $2$ and $\langle a,a^{g},a^{h}\rangle$ is nilpotent of class at most $3$ for all $g,h\in G$. (This is equivalent to $a$ being left $3$-Engel when $a$ is of odd order.) It is proved that if $a$ is a strong left $3$-Engel element in an arbitrary group $G$, then $\langle a\rangle^{G}$ is locally nilpotent (A. Hadjievangelou, G. Traustason, Proc. Amer. Math. Soc., 152, no. 4 (2024), 1467–1477). G. Traustason
@@
If it can, then the answer to 20.117 is negative, but in the abstract case it is known that it can not (P. H. Kropholler, M. A. Roller, J. London Math. Soc. (2), 39 (1989), 271–284). P. Zalesski
@@
Yes, it is true (T. C. Burness, H. Y. Huang, Preprint, 2025, https://arxiv.org/pdf/2506.19745).
@@
b) Yes, it is true (I. N. Belousov, V. I. Zenkov, Preprint, 2025 (Russian),
https://kourovkanotebookorg.wordpress.com/wp-content/uploads/2025/09/20.123b.pdf)
c) Yes, it is true (I. N. Belousov, V. I. Zenkov, Trudy Inst. Mat. Mekh. Ural. Otdel. Ross. Akad. Nauk, 31, no. 1 (2025), 19–35 (Russian)).
@@
Yes, such groups do exist (V. S. Guba, Math. USSR–Izv., 29 (1987), 233–277).
@@
Yes, such groups do exist (S. D. Brodskiĭ, Dep. no. 2214-80, VINITI, Moscow, 1980 (Russian)).
@@
Yes, such groups do exist (L. A. Bokut’, Soviet Math. Dokl., 8 (1967), 901–904; A. Bowtell, J. Algebra, 7 (1967), 126–139; A. Klein, J. Algebra, 7 (1967), 100–125).
@@
Yes, such groups do exist (G. Higman, Finitely presented infinite simple groups ((Notes on Pure Mathematics, no. $8)$ Dept. of Math., Inst. Adv. Studies Austral. National Univ., Canberra, 1974), where there is also a reference to R. Thompson.
@@
a) Yes, there does (G. S. Makanin, Math. USSR–Izv., 21 (1982), 546–582; 25 (1985), 75–88).
@@ 1.45. Is the largest locally nilpotent normal subgroup relatively convex?
Yes, such groups do exist (V. V. Bludov, Algebra and Logic, 13 (1974), 343–360).
@@ 1.61. Can any orderable group be embedded into an orderable group
a) Yes, such groups do exist; b) Not necessarily (S. I. Adian, Math. USSR–Izv., 5 (1971), 475–484).
@@
Yes, such groups do exist (E. M. Levich, A. I. Tokarenko, Siberian Math. J., 11 (1970), 1033–1034; A. I. Tokarenko, Trudy Rizhsk. Algebr. Sem., Riga Univ., Riga, 1969, 280–281 (Russian)).
@@ 1.76. Does there exist a simple, locally nilpotent, locally compact, topological group?
Yes, such groups do exist (V. S. Atabekyan, Dep. no. 5381-V86, VINITI, Moscow, 1986 (Russian)).
@@ 1.85. Is it true that the identity relations of a metabelian group have a finite basis?
Yes, there do (S. V. Ivanov, Trans. Moscow Math. Soc., 1993, 217–249).
@@
Yes, such groups do exist (S. I. Adian, Proc. Steklov Inst. Math., 112 (1971), 61–69).